OpenStudy (anonymous):

Tickets sold in advance for a play cost $3 each. Tickets sold at th edoor for the same play cost $5 each. If 100 tickets were sold for this play for a total of $460, how many tickets were purchased at the door?

OpenStudy (anonymous):

Let the tickets sold in advance be x And let the tickets sold at the edoor be y. So, according to the question: x + y = 100 (Because total 100 tickets were sold) 3x + 5y = 460 (Because total 100 tickets sold for $460) Multiply first equation by 3, 3x + 3y = 300 ----------------1 3x + 5y = 460 -----------------2 Subtract them: 3y - 5y = 300 - 460 -2y = 160 2y = 160 Now can you find y???

OpenStudy (anonymous):

where did you get 3x+3y=300?

OpenStudy (anonymous):

I have just multiply 3 by my own so that the solving method becomes easy... There is one more method to solve it.. Would you solve it on your own?? I will tell you the steps???

OpenStudy (anonymous):

sure, let's try it.

OpenStudy (anonymous):

Firstly, you should about this method.. This method is called SUBSTITUTION METHOD.. Now tell me what is the first equation???

OpenStudy (anonymous):

x+y=100?

OpenStudy (anonymous):

Yes, now find x from here.. can you do that??

OpenStudy (anonymous):

times it by 3, so it will be 3x+3y=300?

OpenStudy (anonymous):

No, i am saying from this equation can you eliminate x, ie x + y = 100 What if I subtract y from both sides, then what will you get??

OpenStudy (anonymous):

like x+y=100 -x+y=100 y=100

OpenStudy (anonymous):

wait, y=-x+100?

OpenStudy (anonymous):

Good but here you have eliminated y.. But you can eliminate y also, there is nothing wrong in that... y = -x + 100 S0, Can I write it as: y = 100 -x. I have just reversed that nothing more.. Right???

OpenStudy (anonymous):

well its minues 100 now, not Plus

OpenStudy (anonymous):

100 is positive and x is negative.. See it with microscope.. Seen or not?? ha ha ha..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So, put this value now in second equation: what is the second equation??

OpenStudy (anonymous):

3x+5y=460 the orginainal equation

OpenStudy (anonymous):

Now in place of y put (100 - x) can you do it or not??

OpenStudy (anonymous):

im trying to figure out what your saying

OpenStudy (anonymous):

See like this: 3x + 5(100 - x) = 460 3x + 500 - 5x = 460 -2x = 460 - 500 2x = 40 x = 20 And now y = 100 - x = 100 - 20, subtract it and get the answer...

OpenStudy (anonymous):

oh okay,t hank you!

OpenStudy (anonymous):

Welcome dear..